OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. x - 4y = 1 2x - 9y = 3 (-1, -3) (-3, 1) (3, 1) (-3, -1)

OpenStudy (anonymous):

x= 1 +4y x= (3 + 9y)/2 (3 + 9y)/2 = 1 +4y 3 + 9y = 2 + 8y y = -1 Then plug that in back two one of those equations.

OpenStudy (anonymous):

No point in just giving you the answer if they are asking you to use the substitution method

OpenStudy (anonymous):

i dont need to show my work just the answer